OpenStudy (anonymous):

1)Verify the identity. Show your work. cot2x + csc2x = 2csc2x - 1  cot²x + csc²x = 2csc²x − 1   cos²x   1      2  = ▬▬▬ + ▬▬▬ = ▬▬▬ − 1    sin²x      sin²x     sin²x 2)Verify the identity. Show your work. 1 + sec2xsin2x = sec2x my work plzz check if wrong plz correct and explain os^2x+sin^2x=1 so (if you divide both sides by cos^2x)... 1+tan^2x=sec^2x tan^2x=sec^2x-1 sec^2x/sec^2x-1=csc^2x sec^2x/tan^2x=csc^2x (1/cos^2x)/(sin^2x/cos^2x)=csc^2x (1/cos^2x)(cos^2x/sin^2x)=csc^2x (1/sin^2x)=csc^2x csc^2x=csc^2x

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathmath333 here it is

OpenStudy (anonymous):

for Q-1 Nnesha heped me but i am cofused if it was solved?

OpenStudy (anonymous):

cot=cos/sin sin=1/csc csc=3 cos=1/sec sec=1/cos is it okay?

OpenStudy (mathmath333):

1st question is this \[\cot2x + \csc2x = 2\csc2x - 1\] or this \(cot^2x + csc^2x = 2csc^2x - 1\)

OpenStudy (mathmath333):

u should only post question in the box that extra thing is confusing me

OpenStudy (anonymous):

oh i am sorry second one is the q u wrote

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ cot^2x + cosec^2x \\~\\ =cosec^2x+cot^2x \\~\\ =2cosec^2x-cosec^2x+cot^2x \\~\\ =2cosec^2x-\dfrac{1}{sin^2x}+\dfrac{cos^2x}{sin^2x} \\~\\ =2cosec^2x+\dfrac{-1+cos^2x}{sin^2x} \\~\\ =2cosec^2x+\dfrac{-1+cos^2x}{sin^2x} \\~\\ =2cosec^2x+\dfrac{-sin^2x}{sin^2x} \\~\\ =2cosec^2x-1 \\~\\ }\end{align}\)

OpenStudy (anonymous):

thanks can post few more? @mathmath333

OpenStudy (mathmath333):

\[\huge 1 + \sec2xsin2x = \sec2x \\or~this\\ \huge1 + \sec^2xsin^2x = \sec^2x \]

OpenStudy (anonymous):

second one @mathmath333 (^) this means raise the power of